OpenStudy (anonymous):

please helpppp Kelly invested $1875 in a savings account which earns 3.35% interest compounded quarterly. What will the account be worth in 7 years? $3,292.34 $1,987.72 $4,717.31 $2,368.20

OpenStudy (anonymous):

@NaCl

OpenStudy (anonymous):

Use this M=P(1+(i)/(n))^(n*t)

OpenStudy (anonymous):

M=1875(1+(0.0335)/(4))^(4*7) = ?

OpenStudy (anonymous):

how do i do ^ on a calculator

OpenStudy (anonymous):

lol you dont !

OpenStudy (anonymous):

copy n paste it onto here http://www.wolframalpha.com/

OpenStudy (anonymous):

and tell me what you get @kenbaby

OpenStudy (anonymous):

@NaCl

OpenStudy (anonymous):

2,368.20 is your answer

OpenStudy (anonymous):

I got that too :) thank you!

OpenStudy (anonymous):

yw :D